# OccamVTS: Distilling Vision Models to 1% Parameters for Time Series Forecasting

### v1

@@ 4.5 Short-term Forecasting
To evaluate performance on short-term prediction tasks, we conduct experiments with forecasting horizons ranging from 6 to 48 time steps across multiple datasets. As shown in Table 4, our method consistently demonstrates 1.3% improvement in SMAPE compared to traditional approaches and maintains significant consistent advantages over the variant without knowledge distillation, with improvements of 1.2-1.5% across all metrics. This validates the effectiveness of our knowledge distillation mechanism even in short-term forecasting scenarios, where the distilled visual features help capture subtle fine-grained temporal patterns that might be missed by pure temporal models. The results suggest that OccamVTS maintains remarkably robust temporal modeling capabilities across different prediction horizons, from short-term operational forecasting to long-term strategic planning.
@@ 4.6 Model Analysis
Ablation Study. To validate the effectiveness of each component, we first conduct ablation experiments on the undistilled teacher model. As shown in Table 5, removing the vision component causes a 6.7% MSE degradation on the Weather dataset, confirming that visual features meaningfully enhance temporal modeling. The more severe 39.7% MSE increase when removing temporal components validates that both modalities are essential, with visual augmentation complementing core temporal representations.
@@
| Avg | 0.224 | 0.259 | 0.239 | 0.276 | 0.313 | 0.343 |
@@ Appendix A Dataset Details
| * The forecastability is calculated by one minus the entropy of Fourier decomposition of time series (Goerg 2013). A larger value indicates better predictability. | ||||||||
@@ Appendix B Evaluation Metrics
For evaluation purposes, we employ mean squared error (MSE) and mean absolute error (MAE) as metrics for long-term forecasting tasks. In contrast, short-term forecasting performance on the M4 benchmark is assessed using symmetric mean absolute percentage error (SMAPE), mean absolute scaled error (MASE), and overall weighted average (OWA), adhering to the evaluation framework established by N-BEATS (Oreshkin et al. 2019). Note that OWA represents a specialized metric designed for the M4 competition. The computation of these metrics follows the formulations below:
@@ 2 Training Parameters
To ensure fair comparison across all methods, we adopt identical experimental settings: 70%/10%/20% train/validation/test splits, consistent hardware infrastructure, and standardized evaluation metrics (MSE and MAE). We embrace a comprehensive training methodology encompassing both general and task-specific parameter configurations, as summarized in Table 8. The model undergoes training with a batch size of and an initial learning rate of , employing the AdamW optimizer (Loshchilov et al. 2017). Early stopping mechanisms are implemented with a patience threshold of epochs to circumvent overfitting. The training process adopts Mean Squared Error (MSE) as the principal loss function and executes for a maximum of epochs. For temporal sequence processing, we utilize an input sequence length of , with prediction horizons configured to , , , or according to task specifications. The output dimension () varies across datasets: for ETTh1/h2/m1/m2, for Weather, for Electricity, and for Traffic. Periodicity parameters are meticulously tailored to dataset characteristics: for ETTh1/h2, Electricity, and Traffic; for ETTm1/m2; and for Weather, ensuring harmonious alignment with dataset-specific temporal patterns. A normalization coefficient of is applied to stabilize training dynamics. The patch embedding module employs a patch length of , stride of , and padding of to process input sequences. The temporal memory mechanism harnesses learnable queries and attention heads to capture sophisticated high-level dependencies.
@@ 4 Teacher and Student Model
The knowledge distillation architecture supports six heterogeneous teacher models—CLIP (512-dimensional), MAE-Base (768-dimensional), MAE-Large (1024-dimensional), MAE-Huge (1280-dimensional) (He et al. 2022), EfficientNet-B3 (1536-dimensional) (Tan et al. 2019), and ResNet101 (2048-dimensional) (He et al. 2016)—while all student models share a 128-dimensional hidden size and adopt lightweight backbones such as Tiny-ViT, EfficientNet-B0, and MobileNet-V3 (Howard et al. 2017). All visual baselines (teachers and students alike) are reproduced from their official publicly released pre-trained weights to ensure fair comparisons.
@@ Appendix D Long-term Forecasting
| Methods | Ours | Only Teacher | Only Student | Time-VLM | TimeMixer++ | TimeMixer | LDM4TS | TimesNet | iTransformer | DLinear | PatchTST | FEDformer | Autoformer | ||||||||||||||
@@ Appendix E Few-shot Forecasting
| Methods | Ours | Only Teacher | Only Student | Times-VLM | TimeMixer++ | TimeMixer | LDM4TS | TimesNet | iTransformer | DLinear | PatchTST | FEDformer | Autoformer | ||||||||||||||
@@
| Methods | Ours | Only Teacher | Only Student | Times-VLM | TimeMixer++ | TimesMixer | LDM4TS | TimesNet | iTransformer | DLinear | PatchTST | Autoformer | |||||||||||||

### v2

@@ 4.5 Short-term Forecasting
To evaluate performance on short-term prediction tasks, we conduct experiments with forecasting horizons ranging from 6 to 48 time steps across multiple datasets. As shown in Table 4, our method consistently demonstrates 1.3% improvement in SMAPE compared to traditional approaches and maintains significant consistent advantages over the variant without knowledge distillation, with improvements of 1.4-2% across all metrics. This validates the effectiveness of our knowledge distillation mechanism even in short-term forecasting scenarios, where the distilled visual features help capture subtle fine-grained temporal patterns that might be missed by pure temporal models. The results suggest that OccamVTS maintains remarkably robust temporal modeling capabilities across different prediction horizons, from short-term operational forecasting to long-term strategic planning.
@@ 4.6 Model Analysis
| 96 | 0.150 | 0.201 | 0.172 | 0.225 | 0.269 | 0.314 |
| 192 | 0.197 | 0.245 | 0.210 | 0.255 | 0.293 | 0.330 |
| 336 | 0.248 | 0.285 | 0.252 | 0.287 | 0.322 | 0.348 |
| 720 | 0.323 | 0.343 | 0.323 | 0.338 | 0.367 | 0.378 |
| Avg | 0.229 | 0.268 | 0.239 | 0.276 | 0.313 | 0.343 |
Ablation Study. To validate the effectiveness of each component, we first conduct ablation experiments on the undistilled teacher model. As shown in Table 5, removing the vision component causes a 4.4% MSE degradation on the Weather dataset, confirming that visual features meaningfully enhance temporal modeling. The more severe 36.7% MSE increase when removing temporal components validates that both modalities are essential, with visual augmentation complementing core temporal representations.
@@ 5 Conclusion
Acknowledgments
This work was supported by the Guangdong Basic and Applied Basic Research Foundation (Grant Nos. 2025A1515011994, 2023B1515120057). It was also supported by the National Key R&D Program of China (Grant No. 2023YFF0725001); the National Natural Science Foundation of China (Grant Nos. 62402414, 92370204); the Guangzhou Municipal Science and Technology Project (Grant No. 2023A03J0011); the Guangzhou Industrial Information and Intelligent Key Laboratory Project (Grant No. 2024A03J0628); a grant from the State Key Laboratory of Resources and Environmental Information System; the Guangdong Provincial Key Laboratory of Integrated Communication, Sensing and Computation for Ubiquitous Internet of Things (Grant No. 2023B1212010007); the Key‑Area Special Project of Guangdong Provincial Ordinary Universities (Grant No. 2024ZDZX1007); and the Education Bureau of Guangzhou, which we gratefully acknowledge.
@@ Appendix A Dataset Details
| * The forecastability is calculated by one minus the entropy of Fourier decomposition of time series (goerg2013forecastable). A larger value indicates better predictability. | ||||||||
@@ Appendix B Evaluation Metrics
For evaluation purposes, we employ mean squared error (MSE) and mean absolute error (MAE) as metrics for long-term forecasting tasks. In contrast, short-term forecasting performance on the M4 benchmark is assessed using symmetric mean absolute percentage error (SMAPE), mean absolute scaled error (MASE), and overall weighted average (OWA), adhering to the evaluation framework established by N-BEATS (oreshkin2019n). Note that OWA represents a specialized metric designed for the M4 competition. The computation of these metrics follows the formulations below:
@@ 2 Training Parameters
To ensure fair comparison across all methods, we adopt identical experimental settings: 70%/10%/20% train/validation/test splits, consistent hardware infrastructure, and standardized evaluation metrics (MSE and MAE). We embrace a comprehensive training methodology encompassing both general and task-specific parameter configurations, as summarized in Table 8. The model undergoes training with a batch size of and an initial learning rate of , employing the AdamW optimizer (loshchilov2017decoupled). Early stopping mechanisms are implemented with a patience threshold of epochs to circumvent overfitting. The training process adopts Mean Squared Error (MSE) as the principal loss function and executes for a maximum of epochs. For temporal sequence processing, we utilize an input sequence length of , with prediction horizons configured to , , , or according to task specifications. The output dimension () varies across datasets: for ETTh1/h2/m1/m2, for Weather, for Electricity, and for Traffic. Periodicity parameters are meticulously tailored to dataset characteristics: for ETTh1/h2, Electricity, and Traffic; for ETTm1/m2; and for Weather, ensuring harmonious alignment with dataset-specific temporal patterns. A normalization coefficient of is applied to stabilize training dynamics. The patch embedding module employs a patch length of , stride of , and padding of to process input sequences. The temporal memory mechanism harnesses learnable queries and attention heads to capture sophisticated high-level dependencies.
@@ 4 Teacher and Student Model
The knowledge distillation architecture supports six heterogeneous teacher models—CLIP (512-dimensional), MAE-Base (768-dimensional), MAE-Large (1024-dimensional), MAE-Huge (1280-dimensional) (he2022masked), EfficientNet-B3 (1536-dimensional) (tan2019efficientnet), and ResNet101 (2048-dimensional) (he2016deep)—while all student models share a 128-dimensional hidden size and adopt lightweight backbones such as Tiny-ViT, EfficientNet-B0, and MobileNet-V3 (howard2017mobilenets). All visual baselines (teachers and students alike) are reproduced from their official publicly released pre-trained weights to ensure fair comparisons.
@@ Appendix D Long-term Forecasting
| Methods | Ours | Only Teacher | Only Student | TimeVLM | TimeMixer++ | TimeMixer | LDM4TS | TimesNet | iTransformer | DLinear | PatchTST | FEDformer | Autoformer | ||||||||||||||
@@ Appendix E Few-shot Forecasting
| Methods | Ours | Only Teacher | Only Student | TimeVLM | TimeMixer++ | TimeMixer | LDM4TS | TimesNet | iTransformer | DLinear | PatchTST | FEDformer | Autoformer | ||||||||||||||
@@
| Methods | Ours | Only Teacher | Only Student | TimeVLM | TimeMixer++ | TimesMixer | LDM4TS | TimesNet | iTransformer | DLinear | PatchTST | Autoformer | |||||||||||||